OpenStudy (dustinrathke24):

A city lot has the shape of a right triangle whose hypotenuse is 3 ft longer than one of the other sides. The perimeter of the lot is 396 ft. How long is each side of the lot?

OpenStudy (jdoe0001):

have you covered quadratic equations yet? like, factoring quadratics by using FOIL

OpenStudy (dustinrathke24):

Yes

OpenStudy (wolf1728):

I made a graphic. That should help.

OpenStudy (dustinrathke24):



OpenStudy (dustinrathke24):

What do I do with this?

OpenStudy (wolf1728):

y^2 = (x+3)^2 -x^2 That should be y^2 = (x+3)^2 -x^2 NOT + x^2 I'd say multiply it out

OpenStudy (wolf1728):

y^2 = (x+3)^2 -x^2 y^2 = x^2 + 6x + 9 -x^2 y^2 = 6x +9 y = sq root (6x + 9)

OpenStudy (dustinrathke24):

Okay now I have \[y^2=6x+9\]

OpenStudy (dustinrathke24):

Okay \[y=\sqrt{6x+9}\]

OpenStudy (wolf1728):

perimeter = 396 So, 396 = (x+3) + x + sq root (6x +9)

OpenStudy (wolf1728):

Square both sides 396^2 = x^2 + 6x + 9 + x^2 + 6x +9

OpenStudy (wolf1728):

156,816 = 2x^2 +12x +18 2x^2 +12x -156,798 =0 Can you solve that quadratic equation?

OpenStudy (dustinrathke24):

yeah

OpenStudy (wolf1728):

I'm thinking I didn't square that correctly. 396 = (x+3) + x + sq root (6x +9) 396^2 = (2x + 3 +sq root (6x+9))^2

OpenStudy (wolf1728):

(2x + 3 +sq root (6x+9))^2 = (2x + 3 + sq root(6x +9) * (2x + 3 + sq root(6x +9)

OpenStudy (wolf1728):

Even if that is squared properly, there will still be a square root radical left in there.

OpenStudy (mathmale):

The side lengths are not necessarily integers, so if one or more length comes out to contain a radical, don't worry about it.

OpenStudy (mathmale):

Summing up the 3 side lengths: x + y + (x+3). Simplify this expression, and then set it = to 396. Solve the resulting equation for y. Then, substitute your expression for y into your diagram: Replace y with it. Solve the resulting equation for x. Finally, find y.